6 + 5 times 2 - 3 + 8 divided 2 times 3 + 1 =26
Answer by stanbon(75887)

6 + 5 times 2 - 3 + 8 divided 2 times 3 + 1 =26
----------------
6 + 5*2 -3 + 8/2 *3 + 1 = 26
----
6 + 10 -3 + 4*3+1 = 26
---
13 +12+1 = 26
26 = 26
===============
Cheers,
Stan H.
===============